### v1

# Tokenizing Buildings: A Transformer for Layout Synthesis

Abstract
We introduce Small Building Model (SBM), a Transformer-based architecture for layout synthesis in Building Information Modeling (BIM) scenes. We address the question of how to tokenize buildings by unifying heterogeneous feature sets of architectural elements into sequences while preserving compositional structure. Such feature sets are represented as a sparse attribute-feature matrix that captures room properties. We then design a unified embedding module that learns joint representations of categorical and possibly correlated continuous feature groups. Lastly, we train a single Transformer backbone in two modes: an encoder-only pathway that yields high-fidelity room embeddings, and an encoder–decoder pipeline for autoregressive prediction of room entities—referred to as Data-Driven Entity Prediction (DDEP). Experiments across retrieval and generative layout synthesis show that SBM learns compact room embeddings that reliably cluster by type and topology, enabling strong semantic retrieval. In DDEP mode, SBM produces functionally sound layouts—with fewer collisions and boundary violations and improved navigability.
Figure 1: Small Building Model (SBM) is an encoder-decoder Transformer that generates functionally correct and semantically coherent layouts given a room envelope. Each row shows a different room type. Our approach outperforms general-purpose LLMs/VLMs and domain-specific methods.
1 Introduction
Professional Computer Aided Design (CAD)/BIM scenes encode rich semantics, hierarchies, and domain constraints, but most strong 3D generators—voxel, mesh, point-cloud, or image-conditioned diffusion models—treat scenes as unstructured geometry [eastman_spatial_1975], yielding visually plausible yet hard-to-edit outputs that frequently violate basic validity rules. Room-scale layout design further demands multi-scale reasoning and parametric outputs compatible with downstream BIM tools, reflecting the highly constrained yet repetitive nature of practice, where designers position walls, doors, casework, and circulation to satisfy codes and programs across hundreds of similar rooms [alphabet_carpo, carpo2017second, eastman_use_1975, Flager2007ACO]. Decades of computational approaches—from CAD macros and parametric families to rule systems, optimization, and learning-based methods—have aimed to automate this process [MONEDERO2000369], but progress remains bottlenecked not by algorithms themselves but by the representations they operate on. An effective representation must expose structure, generalize across building types, and maintain BIM-level geometric and semantic coherence, underscoring the need to rethink spatial data encoding as the foundation for scalable layout reasoning.
To this end, we introduce a normalized hierarchical tokenizer that encodes room topology, entity attributes, wall-referenced geometry, and relational structure into compact BIM-Token Bundles. We realize these bundles as rows of a sparse attribute–feature matrix and embedded via a mixed-type module that produces a unified token vector for Transformer models. Building on this representation, our SBM provides a Transformer backbone for BIM scenes, supporting both encoder-only analysis of room embeddings at scale and encoder–decoder generation through DDEP, which autoregressively places room-level entities with mixed categorical and continuous attributes while preserving the compositional structure of BIM data.
In summary, we contribute (1) a normalized hierarchical tokenization of BIM scenes with a mixed-type embedding module, (2) a unified Transformer backbone (SBM) that supports both retrieval and generative layout synthesis through encoder-only and encoder–decoder (DDEP) modes, and (3) a comprehensive evaluation on a professional BIM dataset demonstrating consistently higher scene coherence, geometric validity, and constraint satisfaction than recent LLM/VLM and domain-specific baselines.
Figure 2: Model overview. (a) BIM data extraction and assembly into a discrete set of token bundles. (b) SBM encoder stack processes the tokenized feature-attribute matrix and outputs a room representation. (c) SBM decoder stack consumes the room representation as memory to the cross-attention layers and the room entities as inputs, trained on next token prediction. (d) Use cases: our SBM is used for three main tasks: DDEP, information retrieval, and user-guided DDEP with the help of an agentic layer.
2 Related Work
Rule-based and heuristic scene and layout generation.
Early work on automatic furniture layout relied on explicit design rules and expert heuristics encoded as constraints. Merrell et al. [merrell_interactive_2011] incorporated interior design guidelines into a cost function and generated layout suggestions by sampling from it. Similarly, Song et al. [song_web3d-based_2019] divided layouts into a few usage modes and applied mode-specific optimization strategies such as recursive subdivision and floor-field methods. Beyond these deterministic rule systems, some approaches convert design guidelines into objective terms and use search-based optimization—such as genetic algorithms—to explore a wider solution space while still operating on the same rule set [kan_automated_2017].
Deep learning–based scene and layout generation.
Graph-based generative models represent layouts as scene graphs and synthesize object arrangements by predicting node attributes and relational edges [xu2018graph2seqgraphsequencelearning, hu2020graph2plan]. Image- and feature-based neural models—ranging from GANs and transformers [tanasra_automation_2023, liu_exploration_2024] to hybrid vectorized systems like HouseDiffusion [shabani2022housediffusion]—learn strong statistical priors for furniture placement yet still depend heavily on training distributions and often require post-hoc refinement to satisfy geometric and functional constraints. Diffusion-based models further improve controllability and semantic consistency, via conditional GAN refinement [nauata2021housegan], mixed discrete–continuous diffusion for 3D scenes [midiffusion2024], room-mask-conditioned semantic diffusion [semLayoutDiff2025], multi-view RGB-D generation [nguyen2024housecrafter], and text-to-layout diffusion Transformers [layyourscene2025, directlayout2025].
LLM/VLM-based scene and layout generation.
Language-based models have recently been applied to spatial layout generation by representing scenes as token sequences. LayoutGPT [feng2023layoutgpt] frames layout synthesis as compositional planning, generating object–relation tokens autoregressively, though it remains reliant on prompting and lacks continuous geometric grounding. Vision-language approaches such as LayoutVLM [sun2025layoutvlm] embed spatial cues jointly with textual descriptions to optimize 3D layouts, yet their coupling of vision features and natural language does not yield a normalized geometric token space.
Other systems use LLMs for guided refinement: HouseTune [zong2024housetune] proposes and linguistically adjusts floorplan candidates, while FlairGPT [littlefair2025flairgpt] aligns language tokens with furniture and style attributes for interior design exploration. Multi-modal models like LLM4CAD [li2024llm4cad] tokenize CAD elements symbolically to support reasoning but operate at the object level without encoding continuous spatial attributes or wall-referenced coordinates. Recently, FloorPlan-DeepSeek [yin2025floorplandeepseek] formulates floorplan synthesis as autoregressive next-room prediction, encoding each room as a vector of semantic type and geometric attributes. This sequential approach constructs layouts incrementally from partial plans or text prompts, achieving improved structural coherence and enabling downstream editing. However, the approach remains limited to room-level floor-plan arrangement, and does not encode fine-grained BIM elements or wall-hosted entities.
Gaps.
Spatial layout research has evolved from rule-based systems to graphs, images, vectorized polygons, and most recently language-based token sequences, each step improving expressiveness. Yet existing representations remain rigid, task-specific, or fragmented, and still fail to unify semantic, relational, and geometric information in a sequence-native form. Most methods yield images, semantic maps, or coarse graphs rather than structured tokens with explicit geometric references and mixed categorical–continuous attributes, limiting constraint-aware reasoning, cross-task generalization, and downstream BIM editability.
This persistent representational gap motivates our approach: a normalized tokenization of BIM scenes coupled with a custom embedding module and a shared Transformer backbone, designed to provide a unified, expressive, and geometrically grounded foundation for scalable spatial layout reasoning.
3 Method
3.1 Problem Setup and BIM Room Representation
We operate on room-level layouts extracted from professional BIM projects (see Sec. 4.1.1 for dataset details). Each room is treated as a parametric layout with semantic attributes, and wall, openings, and furniture/casework entities, all expressed in a local coordinate frame normalized for position and scale.
Envelope vs. contents.
We explicitly decompose each room into a room envelope—the walls, openings, and coarse layout attributes—and a set of room contents (props and casework). We write $r_{\text{env}}=\big(y^{\text{topo}},y^{\text{layout}},\mathcal{E},\mathcal{D},\mathcal{W}\big),r_{\text{ent}}=\big(\mathcal{P},\mathcal{C}\big)$ and define the full room as $r=(r_{\text{env}},r_{\text{ent}})$. Here $y^{\text{topo}}$ is a discrete room-type token (e.g., bedroom, kitchen) and $y^{\text{layout}}$ is a layout token that aggregates global scalar attributes such as area and perimeter. The sets $\mathcal{E},\mathcal{D},\mathcal{W}$ are walls, doors, and windows, and $\mathcal{P},\mathcal{C}$ are props and casework, respectively. In our model, the encoder operates only on the envelope $r_{\text{env}}$, while the decoder operates only on the entity sequence $r_{\text{ent}}$ conditioned on the encoder output.
Walls and room topology.
The room boundary and internal partitions are represented as a set of $N_{E}$ directed wall segments $\mathcal{E}=\{e_{j}\}_{j=1}^{N_{E}},e_{j}=\big(x^{(1)}_{j},x^{(2)}_{j},a^{\text{wall}}_{j}\big),$ where $x^{(1)}_{j},x^{(2)}_{j}\in\mathbb{R}^{2}$ are the endpoints in the local coordinate frame and $a^{\text{wall}}_{j}$ collects wall attributes (e.g., inside/outside thickness, wall condition or construction type, and a discrete edge identifier). The counter-clockwise ordered set $\mathcal{E}$ defines the room polygon and its topology; global geometric scalars such as area and perimeter are computed from $\mathcal{E}$ and packed into the layout token $y^{\text{layout}}$.
Openings: doors and windows.
Doors and windows are modeled as objects attached to specific wall edges with wall-referenced coordinates. For doors we write $\mathcal{D}=\{d_{k}\}_{k=1}^{N_{D}},d_{k}=\big(j_{k},t_{k},w_{k},a^{\text{door}}_{k}\big),$ where $j_{k}\in\{1,\dots,N_{E}\}$ indexes the supporting edge $e_{j_{k}}$, $t_{k}\in[0,1]$ is a normalized position along that edge, $w_{k}$ is the opening width, and $a^{\text{door}}_{k}$ includes additional categorical attributes (e.g., door family/type, swing direction). Windows $\mathcal{W}=\{w_{\ell}\}_{\ell=1}^{N_{W}}$ are defined analogously with edge attachment, normalized position, width, and window-specific attributes. This wall-referenced parameterization makes openings invariant to absolute translation and scale and directly compatible with BIM editing operations.
Furniture and casework entities.
Room contents are represented as two sets of entities: $\mathcal{P}=\{p_{u}\}_{u=1}^{N_{P}},\qquad\mathcal{C}=\{c_{v}\}_{v=1}^{N_{C}},$ for props (furniture) and casework. Each entity is parameterized by a discrete type and a small set of wall-referenced continuous attributes. For a generic entity $q\in\mathcal{P}\cup\mathcal{C}$ we write $q=\big(c_{q},j_{q},t_{q},\delta_{q},s_{q},\rho_{q},a^{\text{ent}}_{q}\big),$ where $c_{q}$ is an entity type identifier (e.g., bed, base_cabinet), $j_{q}$ is the supporting edge index, $t_{q}\in[0,1]$ is the position along that edge, $\delta_{q}$ is a signed lateral offset from the wall into the room, $s_{q}$ encodes size or footprint parameters (e.g., width/depth), $\rho_{q}$ is an in-plane rotation angle when applicable (props only), and $a^{\text{ent}}_{q}$ contains any additional categorical attributes. This representation captures both semantic identity and precise placement of entities relative to the room envelope.
Tasks.
Given this decomposition, our model supports two tasks. In encoder-only mode, the goal is to map the room envelope $r_{\text{env}}$ to a compact embedding $z(r_{\text{env}})\in\mathbb{R}^{d}$ that preserves semantic and geometric relationships between rooms (for retrieval and analysis). In encoder–decoder mode (DDEP), the encoder consumes only the envelope $r_{\text{env}}$ to produce a contextual memory, and the decoder autoregressively generates the entity sequence $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ as a sequence of tokens (see Fig. 2). The following subsections describe how we convert $r_{\text{env}}$ and $r_{\text{ent}}$ into BIM-Token Bundles on sparse attribute–feature matrices, embed these mixed-type features, and train a shared Transformer backbone to solve both tasks.
3.2 BIM Tokenization
Given the room envelope $r_{\text{env}}$ and contents $r_{\text{ent}}$ (Sec. 3.1), we convert each room into two ordered sequences of BIM-Token Bundles: an envelope sequence for the encoder and an entity sequence for the decoder. Each bundle corresponds to a single logical element (topology, layout, wall, door, window, or entity) and aggregates all of its semantic and geometric attributes into a sparse feature vector.
Token sequences.
On the encoder side, we construct a fixed-order sequence that starts with a special classification token and ends with an end-of-sequence token:
$\big(\tau^{\text{CLS}},\tau^{\text{t}},\tau^{\text{l}},\tau^{\text{e}}_{1},\dots,\tau^{\text{e}}_{N_{E}},\tau^{\text{d}}_{1},\dots,\tau^{\text{d}}_{N_{D}},\tau^{\text{w}}_{1},\dots,\tau^{\text{w}}_{N_{W}},\tau^{\text{EOS}}\big),$
followed by padding tokens up to a maximum length $S_{\text{enc}}$. The CLS bundle $\tau^{\text{CLS}}$ is a learnable summary token used for pooling in encoder-only mode; it does not correspond to any specific BIM element. The topology bundle $\tau^{\text{t}}$ encodes $y^{\text{topo}}$ (room type), the layout bundle $\tau^{\text{l}}$ encodes $y^{\text{layout}}$ (area, perimeter, and other global scalars), each edge bundle $\tau^{\text{e}}_{j}$ aggregates the attributes of $e_{j}\in\mathcal{E}$, and each door or window bundle aggregates the corresponding element in $\mathcal{D}$ or $\mathcal{W}$ with wall-referenced parameters. The EOS bundle $\tau^{\text{EOS}}$ marks the end of the envelope sequence. This encoder sequence is envelope-only: props and casework are not encoded in the encoder input.
On the decoder side, we form a contents-only entity sequence $\big(\tau^{\text{SOS}},\tau^{\text{ent}}_{1},\dots,\tau^{\text{ent}}_{T},\tau^{\text{EOS}}\big),$ where each $\tau^{\text{ent}}_{t}$ corresponds to one entity $q\in\mathcal{P}\cup\mathcal{C}$ with its type and wall-referenced attributes, and $\tau^{\text{SOS}}$, $\tau^{\text{EOS}}$ are special start/end tokens. The room envelope is provided to the decoder only via cross-attention to the encoder memory, not re-encoded as decoder tokens.
Attribute–feature matrices.
We realize these heterogeneous token sequences as attribute–feature matrices with a fixed feature axis and variable sequence length. For the encoder, we define $X^{\text{enc}}\in\mathbb{R}^{F_{\text{enc}}\times S_{\text{enc}}}$,
$X^{\text{enc}}=\begin{bmatrix}x^{\text{enc}}_{1,1}&x^{\text{enc}}_{1,2}&\cdots&x^{\text{enc}}_{1,S_{\text{enc}}}\\ x^{\text{enc}}_{2,1}&x^{\text{enc}}_{2,2}&\cdots&x^{\text{enc}}_{2,S_{\text{enc}}}\\ \vdots&\vdots&\ddots&\vdots\\ x^{\text{enc}}_{F_{\text{enc}},1}&x^{\text{enc}}_{F_{\text{enc}},2}&\cdots&x^{\text{enc}}_{F_{\text{enc}},S_{\text{enc}}}\end{bmatrix}$,
where each column $X^{\text{enc}}_{:,s}$ encodes one BIM-Token Bundle at sequence position $s$, and each row corresponds to a specific feature. The first rows are two generic categorical identifiers that are present for every token, a type id to denote the semantic attribute, and a token id to index each semantic instance within each type. The remaining rows collect room-level scalars (e.g., area, perimeter), wall geometry (edge endpoints, lengths, thicknesses), and opening parameters (normalized position along the wall, width, corner distances), as configured in a fixed feature table. Only a subset of these features is active for any given token type; non-applicable entries are filled with a designated padding value, yielding a sparse representation. A complete list of encoder feature rows is provided in the supplement.
We construct the decoder matrix $X^{\text{dec}}\in\mathbb{R}^{F_{\text{dec}}\times S_{\text{dec}}}$ in the same way, now with decoder-specific features such as entity type, edge attachment, wall-referenced coordinates $(t_{q},\delta_{q})$, size parameters $s_{q}$, and rotation $\rho_{q}$, plus token_type_id and token_id. Here too, only the rows relevant to each token type are active and the rest are set to p=100p=-100. In the next subsection, we describe how these sparse matrices $X^{\text{enc}}$ and $X^{\text{dec}}$ are mapped to dense token embeddings by a mixed-type embedding module.
3.3 Mixed-Type Embedding Module
The BIM-Token Bundles defined above combine heterogeneous categorical and continuous attributes on a shared feature axis. Our mixed-type embedding module converts the sparse feature matrices $X^{\text{enc}}$ and $X^{\text{dec}}$ into dense token embeddings suitable for a Transformer, while respecting which features are active for which tokens.
Formally, we define two embedding maps
$\Phi_{\text{enc}}:\mathbb{R}^{F_{\text{enc}}\times S_{\text{enc}}}\rightarrow\mathbb{R}^{S_{\text{enc}}\times d},\!\!\!\!\!\!\qquad\Phi_{\text{dec}}:\mathbb{R}^{F_{\text{dec}}\times S_{\text{dec}}}\rightarrow\mathbb{R}^{S_{\text{dec}}\times d},$
implemented by a shared FeatureEmbedding module configured with encoder- or decoder-specific feature definitions. In both cases, each token embedding is constructed by summing the contributions of all its active features.
Feature-wise embeddings and masking.
Consider first the encoder. For a single room, the encoder feature matrix $X^{\text{enc}}$ has rows $x_{f}\in\mathbb{R}^{S_{\text{enc}}}$ corresponding to individual features (e.g., token_type_id, area, edge_x1, t_value). For each feature $f$ we instantiate a feature-specific embedding function $E_{f}$:
  • Categorical features (e.g., token_type_id, token_id, edge indices, wall condition) use learnable embedding tables $E_{f}:\mathbb{Z}\to\mathbb{R}^{d}$ with a dedicated entry for padding.
  • Scalar continuous features (e.g., area, perimeter, edge length, relative length) use small multilayer perceptrons $E_{f}:\mathbb{R}\to\mathbb{R}^{d}$.
  • Grouped continuous features (e.g., edge endpoint coordinates, thickness pairs, door/window corner distances) use a MultiContinuousEmbedding: a lightweight sub-network that embeds each scalar in the group, applies self-attention or pooling across the group, and projects the result to $\mathbb{R}^{d}$.
Let $\mathcal{F}_{\text{enc}}$ be the set of encoder features. For each feature $f\in\mathcal{F}_{\text{enc}}$ and token position $s$, we define a valid-feature mask $m_{f,s}=\mathds{1}\big[X^{\text{enc}}_{f,s}\neq-100\,],$ which indicates whether feature $f$ applies to token $s$. After replacing sentinel values 100-100 by 0 for numerical stability inside $E_{f}$, we compute feature embeddings and apply the mask: $u_{f,s}=E_{f}\big(X^{\text{enc}}_{f,s}\big)\in\mathbb{R}^{d},\tilde{u}_{f,s}=m_{f,s}\cdot u_{f,s}.$ This second masking step ensures that inactive features contribute exactly zero, including the bias terms of linear layers. The encoder attention mask further zeroes out all embeddings for padded token positions.
Token embedding as feature sum.
The final encoder token embedding at position $s$ is the sum of all active feature embeddings for that token: $e^{\text{enc}}_{s}=\sum_{f\in\mathcal{F}_{\text{enc}}}\tilde{u}_{f,s}\in\mathbb{R}^{d}.$ Stacking these over sequence positions yields $E^{\text{enc}}=\Phi_{\text{enc}}(X^{\text{enc}})\in\mathbb{R}^{S_{\text{enc}}\times d},$ optionally augmented with learned positional embeddings. This construction is token-type agnostic: topology, layout, walls, doors, and windows are all embedded via the same mechanism, differing only in which feature rows are active for each token.
Decoder embeddings.
The decoder uses the same FeatureEmbedding architecture with a decoder-specific feature set. Given the decoder feature matrix $X^{\text{dec}}$, we instantiate analogous categorical, continuous, and grouped-continuous embedders for features such as token_type_id, token_id, edge attachments, wall-referenced coordinates $(t_{q},\delta_{q})$, size parameters $s_{q}$, and rotations $\rho_{q}$. Applying the same masking and summation procedure yields decoder token embeddings $E^{\text{dec}}=\Phi_{\text{dec}}(X^{\text{dec}})\in\mathbb{R}^{S_{\text{dec}}\times d},$ with the property that entities of different types (props vs. casework, SOS/EOS vs. regular tokens) share the same embedding space but rely on different subsets of features.
By treating each BIM-Token Bundle as a sparse feature column and mapping it through this mixed-type embedding module, we obtain dense, sequence-native token embeddings for both the room envelope and its contents. The next subsection describes how a standard Transformer backbone consumes $E^{\text{enc}}$ and $E^{\text{dec}}$ in encoder-only and encoder–decoder modes.
3.4 Transformer Backbone and Operating Modes
Given the encoder and decoder token embeddings we employ a standard Transformer encoder–decoder architecture. The same backbone is used in two operating modes: encoder-only (for room embeddings and retrieval) and encoder–decoder (for data-driven entity prediction, DDEP).
Encoder-only: room embeddings and retrieval.
In encoder-only mode, we use only the Transformer encoder. Given $E^{\text{enc}}$ and an attention mask over non-padding tokens, we compute $M=\mathrm{Enc}_{\theta}(E^{\text{enc}})\in\mathbb{R}^{S_{\text{enc}}\times d},$ where $\mathrm{Enc}_{\theta}$ is a stack of self-attention and feed-forward blocks. The first position in the sequence corresponds to the special CLS token $\tau^{\text{CLS}}$ introduced in Sec. 3.2. We obtain a room embedding for the envelope by pooling at this position $z(r_{\text{env}})=M_{0}\in\mathbb{R}^{d},$ which we use directly for room retrieval and similarity-based analysis. Encoder heads (e.g., room-type classification and masked-token prediction) operate on $z(r_{\text{env}})$ and on the full memory $M$.
Encoder–decoder: Data-Driven Entity Prediction (DDEP).
In DDEP mode, we use the full encoder–decoder Transformer to generate the room contents $r_{\text{ent}}=(\mathcal{P},\mathcal{C})$ conditioned on the room envelope $r_{\text{env}}$. The encoder processes the envelope as above, producing a memory $M$ that is held fixed during decoding. The decoder takes entity token embeddings $E^{\text{dec}}$ and attends both to the prefix of generated tokens (causal self-attention) and to the encoder memory $M$ (cross-attention), $H=\mathrm{Dec}_{\theta}\big(E^{\text{dec}},M\big)\in\mathbb{R}^{S_{\text{dec}}\times d},$ where $\mathrm{Dec}_{\theta}$ is a stack of masked self-attention, cross-attention, and feed-forward layers. Decoder states in $H$ feed into multi-head prediction layers that produce mixed categorical and continuous outputs for each entity token. We enforce a constrained structured decoding schema which refines the raw predictions.
4 Experiments and Results
We structure this section around our two main parts, DDEP and SBM embeddings, and compare our results to previous approaches. Due to space constraints, we defer training, implementation details and full tokenizer/adapter configurations to the Supplementary Material. In all experiments, we use DDEP in its unguided form: the decoder autoregressively generates entity sequences conditioned only on the room envelope and encoder memory, without retrieval-based conditioning or external agent feedback.
Coverage Navigability OC (%) Inference Latency (s) Method $\uparrow$ mean $\pm$ std $\uparrow$ mean $\pm$ std $\uparrow$ SR (mean $\pm$ std) $\downarrow$ DF (mean $\pm$ std) $\downarrow$ mean $\pm$ std $\downarrow$ mean $\pm$ std Text-based methods Claude Sonnet 4.5 34.1 ± 0.3 2.1 ± 2.7 0.26 ± 0.02 0.69 ± 0.03 33.3 ± 2.5 18.02 ± 12.19 Gemini 2.5 Pro 67.1 ± 1.6 58.0 ± 8.3 0.69 ± 0.07 0.30 ± 0.03 10.2 ± 1.6 42.51 ± 13.25 GPT-5 Medium 56.8 ± 1.5 51.2 ± 8.8 0.63 ± 0.07 0.32 ± 0.04 12.9 ± 1.8 100.86 ± 65.58 Qwen3 60.5 ± 1.4 34.0 ± 0.8 0.50 ± 0.01 0.45 ± 0.02 18.1 ± 1.6 21.27 ± 13.32 VLM methods Claude Sonnet 4.5 44.9 ± 3.4 52.7 ± 1.7 0.64 ± 0.01 0.32 ± 0.03 23.7 ± 2.0 18.71 ± 3.62 Gemini 2.5 Pro 73.2 ± 1.2 54.6 ± 2.2 0.67 ± 0.02 0.34 ± 0.02 11.9 ± 1.9 43.16 ± 13.16 GPT-5 Medium 55.5 ± 1.4 54.3 ± 10.4 0.65 ± 0.08 0.29 ± 0.05 12.3 ± 3.9 164.43 ± 105.88 GPT-5 High 58.3 ± 0.0 57.8 ± 54.0 0.67 ± 0.42 0.27 ± 0.39 12.1 ± 25.5 124.18 ± 78.34 Qwen3 75.9 ± 1.5 48.7 ± 1.0 0.60 ± 0.02 0.33 ± 0.02 12.4 ± 1.6 439.04 ± 377.43 Domain-specific methods LayoutVLM 40.3 ± 59.8 0.55 ± 0.45 0.41 ± 0.45 4.9 ± 11.4 144.99 ± 36.50 FlairGPT 46.6 ± 36.5 28.2 ± 61.0 0.46 ± 0.46 0.51 ± 0.46 5.0 ± 12.9 1133.61 ± 1596.40 DDEP (Ours) 98.2 ± 6.3 82.4 ± 40.7 0.89 ± 0.31 0.18 ± 0.29 3.2 ± 10.2 3.18 ± 0.18
Table 1: Comparison on the held-out set. We report mean $\pm$ std. Higher is better for Coverage, Navigability (mean and SR); lower is better for OC and DF. Best overall (SOTA) also in bold. LayoutVLM requires an initial furniture list as input, so it is excluded from Coverage metric. Latency is reported as mean $\pm$ std per room, measured end-to-end on an AWS EC2 g5.xlarge instance.
4.1 Experimental Setup
4.1.1 Dataset and Setup
We train SBM on a proprietary corpus of residential BIM scenes containing single-family homes with typed rooms (e.g., bathrooms, dining rooms, offices, pantries). Across these constraints, the dataset contains 439,932 training samples over. We construct held-out splits per room type and report all quantitative results on the same test rooms for every method.
Each room is converted to BIM-Token Bundles and then to encoder and decoder sequences as in Sec. 3.2, yielding relatively short sequences compared to text corpora. Over all room types and training regimes, SBM sees on the order of five million tokens in total across encoder and decoder streams, i.e., a few million envelope tokens and a comparable number of entity tokens. Despite this modest scale, the dataset covers a wide range of residential geometries and furniture programs, providing enough diversity to train both the encoder-only embedding pathway and the DDEP decoder for the room types evaluated in Secs. 4.2 and 4.4.
4.1.2 Compared Methods
We evaluate three families of methods on the same held-out rooms. For each room envelope, we construct a serialized text description and, when applicable, a rendered floorplan image. All methods return predicted props and casework, which we map to a unified schema of entity categories and dimensions. For LLM/VLM baselines we use fixed prompts (see Suppl.) and decoding (T=0.7, top‑p=0.9 unless stated). VLMs condition on a 1024×1024 floorplan render (orthographic, 20px/m). Free‑text outputs are parsed via a rule‑based mapper to our inventory. Each method runs 3 seeds; we report mean +/- sd and inference time. For reproducible implementation details, refer to Supplementary Material.
Text LLMs. Claude Sonnet 4.5, Gemini 2.5 Pro, GPT-5, and Qwen3 in text-only mode.
Vision–language models (VLMs). The same base models with image-conditioned variants that receive a floorplan render plus textual context.
Domain-specific methods. LayoutVLM [sun2025layoutvlm], FlairGPT [littlefair2025flairgpt], and our DDEP model built on the SBM encoder–decoder backbone.
4.1.3 Layout Quality Metrics
We report three complementary layout metrics: coverage (inventory satisfaction), navigability (functional access), and overlap & clearance (spatial violations). All metrics use the same room geometry, entity dimensions, and inventory specification for every method. Full implementation details are provided in the Supplement.
Coverage.
Coverage measures how well a predicted layout $\hat{y}$ satisfies the room-type inventory $y$, independent of exact placement. Let $\mathcal{I}$ be required items and $\mathcal{G}$ mutually exclusive alternative groups with non-negative weights $w_{i},w_{g}$:
$\mathrm{Cov}(y,\hat{y})=\frac{\sum_{i\in\mathcal{I}}w_{i}\,s_{i}(\hat{y})\;+\;\sum_{g\in\mathcal{G}}w_{g}\,s_{g}(\hat{y})\;-\;p_{\text{extra}}(\hat{y})}{\sum_{i\in\mathcal{I}}w_{i}\;+\;\sum_{g\in\mathcal{G}}w_{g}}.$
Item scores $s_{i}\!\in[0,1]$ handle missing/overfill counts; group scores $s_{g}\!\in[0,1]$ capture satisfaction of any valid alternative; $p_{\text{extra}}\!\geq\!0$ penalizes extraneous categories.
Navigability.
The navigability metric evaluates the walkability from door-to-target. Let $\mathcal{P}$ denote all door–target pairs, and let $\ell_{t}$ be the shortest collision-free path length for pair $t$ (computed with A*), with $\ell_{t}^{\star}$ the straight-line distance inside the room. We define the success rate and detour factor as $\mathrm{SR}=\frac{1}{|\mathcal{P}|}\sum_{t\in\mathcal{P}}\mathbf{1}[\ell_{t}<\infty],\quad\mathrm{DF}=\frac{1}{|\mathcal{P}_{\text{reach}}|}\sum_{t\in\mathcal{P}_{\text{reach}}}\frac{\ell_{t}}{\ell_{t}^{\star}},$ where $\mathcal{P}_{\text{reach}}=\{t\in\mathcal{P}:\ell_{t}<\infty\}$. For visualization we also report a scalar navigability score $\mathrm{Nav}=100\times\big(\mathrm{SR}-\lambda\,\mathrm{DF}\big)$, with $\lambda$ chosen on a validation set to balance reachability and detour.
Overlap & clearance.
The overlap & clearance metric measures spatial violations using exact polygon geometry. We compute four normalized terms: entity overlap fraction (EOF), global overlap area (GOA), door clearance intrusion (DCI), and wall-bounds violation (WBV). These are combined into a single overlap/clearance score
$\mathrm{OC}=w_{\text{EOF}}\,\mathrm{EOF}+w_{\text{GOA}}\,\mathrm{GOA}+w_{\text{DCI}}\,\mathrm{DCI}+w_{\text{WBV}}\,\mathrm{WBV},$
with fixed weights $w_{\cdot}$ shared for all methods.
4.2 DDEP Quantitative Comparisons
Figure 3: Qualitative comparison of generated layouts across five room types, showing representative results from seven baseline methods and our DDEP model.
On the held-out benchmark (Table 1), we compare layout generators using three metrics. Text LLMs achieve moderate coverage but weak functional quality (under-furnishing, low reachability, higher OC). VLMs improve some axes but exhibit clear trade-offs (gains in coverage often accompany reduced reachability or more collisions). Domain-specific baselines (LayoutVLM, FlairGPT) show lower violation rates yet remain weak on both coverage and navigability under our unified schema. DDEP occupies a distinct regime: near-complete coverage, the highest navigability, and the lowest OC—despite placing substantially more entities than any baseline. No competing method attains high coverage and high navigability simultaneously without incurring larger spatial errors, positioning DDEP as state of the art on this benchmark.
4.3 Qualitative Comparisons
Figure 3 shows example layouts for several room types from DDEP, VLMs, and domain-specific methods. VLM baselines often overfill rooms with extra casework and furniture, blocking circulation and door clearances, especially in living rooms and garages. Domain-specific methods tend to place more appropriate furniture types but still leave narrow or blocked paths around key elements. In contrast, DDEP produces layouts that satisfy the room program while keeping clear, door-connected circulation bands and respecting wall-referenced anchoring, visually matching the trends reported in our quantitative metrics.
4.4 Embedding Evaluation and Encoder Ablations
We evaluate SBM’s encoder-only pathway for room embedding quality through retrieval and clustering benchmarks. As an ablation study, we compare against state-of-the-art text embedding baselines to demonstrate SBM’s encoder capacity and highlight its complementary capabilities for geometric and spatial understanding. We compare against E5-Large-v2 [wang2022e5]: 560M parameters, pre-trained on Common Crawl and Wikipedia, BGE-Large-en-v1.5 [bge_embedding]: 335M parameters, pre-trained on web-scale text data, and GTE-Large-en-v1.5 [li2023towards]: 335M parameters, pre-trained on web-scale text corpora. These baselines are pre-trained on billions of text tokens, representing orders of magnitude more training data than SBM’s domain-specific dataset. For fair comparison, we serialize room data to text format, enabling text embeddings to process the same room information.
Evaluation Metrics
We assess retrieval quality with nDCG@k (graded relevance) and Success@k/MRR/Precision@k, structural consistency with type-constrained nDCG@k and triplet accuracy, entity-level similarity via overlap of props/casework/doors, and global organization with clustering metrics (NMI/ARI/Silhouette) and geometric rank correlation (Spearman/Kendall) against a geometry oracle. Full metric definitions and cross-type retrieval tables are provided in the Supplement; here we focus on type-constrained retrieval and clustering/correlation.
Method nDCG Triplet Entity Sim Clust.(NMI) SBM 62.6 100.0 78.9 0.640 E5-Large-v2 86.7 99.9 78.4 0.371 BGE-Large-en-v1.5 80.1 99.9 77.0 0.358 GTE-Large-en-v1.5 74.3 99.8 77.0 0.335
Table 2: Type-Constrained Retrieval Performance @ k=5
Method NMI ARI Silhouette Spearman SBM 0.640 0.307 0.152 0.145 E5-Large-v2 0.371 -0.019 0.051 0.315 BGE-Large-en-v1.5 0.358 -0.003 0.069 0.235 GTE-Large-en-v1.5 0.335 -0.002 0.067 0.133
Table 3: Clustering and Geometric Correlation Metrics
Table 2 shows that text encoders lead on ranking quality within type (nDCG: E5 86.7, BGE 80.1, GTE 74.3 vs. SBM 62.6), while SBM excels at structure: it attains perfect triplet accuracy (100.0), the highest entity overlap (78.9), and stronger type organization (NMI 0.640). In short, text models rank better among same-type rooms, whereas SBM preserves within-type geometry and entities more faithfully.
Figure 4: UMAP visualization of room embeddings colored by room type category. SBM embeddings (left, NMI: 0.640) exhibit well-separated clusters with distinct boundaries between room types, demonstrating superior geometric and spatial understanding. E5-Large-v2 embeddings (right, NMI: 0.371) show more intermingled clusters with blurred boundaries, indicating weaker room type separation. The 1.7× higher NMI score for LBM reflects its specialization in capturing geometric structure and spatial relationships inherent in building layouts, rather than semantic similarity alone.
As summarized in Table 3, SBM yields markedly better global organization (NMI/ARI/Silhouette: 0.640/0.307/0.152) than text baselines (NMI 0.335–0.371), but aligns less with our geometry oracle in rank correlation (Spearman 0.145 vs. E5’s 0.315). This indicates a tighter type-structured space for SBM (Fig. 4) but stronger oracle-ranked agreement for E5.
5 Conclusion
We introduced the SBM, a Transformer architecture tailored to BIM-like room layouts through BIM-Token Bundles, sparse attribute–feature matrices, and a mixed-type embedding module shared by encoder and decoder. This representation treats topology, walls, openings, and entities as sequence-native tokens with joint categorical and continuous attributes, enabling a single backbone to support both encoder-only room embeddings and an encoder–decoder pipeline (DDEP). On a BIM-grounded benchmark with specification-aware coverage, navigability, and overlap/clearance metrics, DDEP delivers near-complete inventory satisfaction, state-of-the-art navigability, and the lowest violation rates across LLM/VLM and domain-specific baselines. Encoder ablations show that SBM yields geometry- and entity-aware embeddings that complement generic text encoders, trading some raw ranking strength for substantially better within-type structure. Our study suggests that modestly sized, domain-specific sequence models over well-designed BIM tokenizations can outperform much larger general-purpose systems on layout quality.
Limitations.
Advanced vertical constraints like sloped ceilings or stairs/elevation changes, and MEP/service routing are not modeled. The benchmark largely comprises residential rooms from our corpus, so distributional shifts (styles, scales, local codes) may affect external validity.